A 352-N, uniform, 1.68-m bar is suspended horizontally by two vertical cables at each end. Cable A can support a maximum tension of 516.0 N without breaking, and cable B can support up to 434.0 N . You want to place a small weight on this bar

Part A

What is the heaviest weight you can put on without breaking either cable?

Part B

Where should you put this weight?
Answer #1

a) Net force =0 No = 516 + 434–369 = = 581N b) Not torque caround it is zero Nodz 434 x 1-68 – 369* 0.84 a = 729.12 - 309.96